# Offline Learning for Combinatorial Multi-armed Bandits

### v1

@@ Abstract
The combinatorial multi-armed bandit (CMAB) is a fundamental sequential decision-making framework, extensively studied over the past decade. However, existing work primarily focuses on the online setting, overlooking the substantial costs of online interactions and the readily available offline datasets. To overcome these limitations, we introduce Off-CMAB, the first offline learning framework for CMAB. Central to our framework is the combinatorial lower confidence bound (CLCB) algorithm, which combines pessimistic reward estimations with combinatorial solvers. To characterize the quality of offline datasets, we propose two novel data coverage conditions and prove that, under these conditions, CLCB achieves a near-optimal suboptimality gap, matching the theoretical lower bound up to a logarithmic factor. We validate Off-CMAB through practical applications, including learning to rank, large language model (LLM) caching, and social influence maximization, showing its ability to handle nonlinear reward functions, general feedback models, and out-of-distribution action samples that excludes optimal or even feasible actions. Extensive experiments on synthetic and real-world datasets further highlight the superior performance of CLCB.
@@ 1 Introduction
Combinatorial multi-armed bandit (CMAB) is a fundamental sequential decision-making framework, designed to tackle challenges in combinatorial action spaces. Over the past decade, CMAB has been extensively studied [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17], driving advancements in real-world applications like recommendation systems [18, 19, 20, 21], healthcare [22, 23, 24], and cyber-physical systems [25, 26, 27, 28] .
@@
Most success stories of CMAB have emerged within the realm of online CMAB [5, 6, 7, 8, 9, 10, 11, 21, 29, 13], which relies on active data collection through online exploration. While effective in certain scenarios, this framework faces two major limitations. On one hand, online exploration becomes impractical when it incurs prohibitive costs or raises ethical and safety concerns. On the other hand, they neglect offline datasets that are often readily available at little or no cost. For instance, in healthcare systems [30], recommending optimal combinations of medical treatments—such as drugs, surgical procedures, and radiation therapy—requires extreme caution. Experimenting directly on patients is ethically and practically infeasible. Instead, leveraging pre-collected datasets of prior treatments can help to make informed decisions while ensuring patient safety. Similar happens for recommendation systems [31] and autonomous driving [32], offline datasets such as user click histories and human driving logs are ubiquitous. Leveraging these offline datasets can guide learning agents to identify optimal policies while avoiding the significant costs associated with online exploration—such as degrading user experience or risking car accidents.
@@
To address the limitations of online CMAB, we propose the first offline learning framework for CMAB (Off-CMAB), where we leverage a pre-collected dataset consisting of $n$ samples of combinatorial actions and their corresponding feedback data. Our framework handles rewards that are nonlinear functions of the chosen super arms and considers general probabilistic feedback models, supporting a wide range of applications such as learning to rank [33], large language model (LLM) caching [34], and influence maximization [35]. The objective is to identify a combinatorial action that minimizes the suboptimal gap, defined as the reward difference between the optimal action and the identified action. The key challenge of Off-CMAB lies in the absence of access to an online environment, which inherently limits the number of data samples available for each action. Furthermore, this problem becomes even more challenging with the combinatorially large action space, which complicates the search for optimal solutions, and the potential presence of out-of-distribution (OOD) samples, where the dataset may exclude optimal or even feasible actions. To tackle these challenges, this work makes progress in answering the following two open questions:
@@
Practical Applications: We show the practicality of Off-CMAB by fitting real-world problems into our framework and applying CLCB to solve them, including (a) learning to rank, (b) LLM caching, and (c) social influence maximization (IM). For the LLM cache problem, beyond directly fitting it into our framework, we improve existing results by addressing full-feedback arms, extending our approach to the online LLM setting with similar improvements. For social IM, our framework handles nuanced node-level feedback by constructing base-arm LCBs via intermediate UCB/LCBs, with additional refinements using variance-adaptive confidence intervals for improved performance.
@@
1.1 Related Works
1.1.1 Combinatorial Multi-armed Bandits
The combinatorial multi-armed bandit (CMAB) problem has been extensively studied over the past decade, covering domains such as stochastic CMAB [5, 6, 7, 8, 9, 10, 11, 21, 29, 13], adversarial CMAB [25, 36, 1, 2, 3, 4, 37], and hybrid best-of-both-worlds settings [12, 38, 39]. Contextual extensions with linear or nonlinear function approximation have also been explored [14, 40, 15, 41, 42, 16, 17, 13].
Our work falls within the stochastic CMAB domain, first introduced by Gai et al. [5], with a specific focus on CMAB with probabilistically triggered arms (CMAB-T). Chen et al. [8] introduced the concept of arm triggering processes for applications like cascading bandits and influence maximization, proposing the CUCB algorithm with a regret bound of $O(B_{1}\sqrt{mKT\log T}/p_{\min})$ regret bound under the 1-norm smoothness condition with coefficient $B_{1}$. Subsequently, Wang and Chen [9] refined this result, proposed a stronger 1-norm triggering probability modulated (TPM) $B_{1}$ smoothness condition, and employed triggering group analysis to eliminate the $1/p_{\min}$ factor from the previous regret bound. More recently, Liu et al. [29] leveraged the variance-adaptive principle to propose the BCUCB-T algorithm, which further reduces the regret’s dependency on action-size from $O(K)$ to $O(\log K)$ under the new variance and triggering probability modulated (TPVM) condition. While inspired by these works, our study diverges by addressing the offline CMAB setting, where online exploration is unavailable, and the focus is on minimizing the suboptimality gap rather than regret.
1.1.2 Offline Bandit and Reinforcement Learning
Offline reinforcement learning (RL), also known as “batch RL", focuses on learning from pre-collected datasets to make sequential decisions without online exploration. Initially studied in the early 2000s [43, 44, 45], offline RL has gained renewed interest in recent years [46].
From an empirical standpoint, offline RL has achieved impressive results across diverse domains, including robotics [47], healthcare [30], recommendation systems [31], autonomous driving [32], and large language model fine-tuning and alignment [48]. Algorithmically, offline RL approaches can be broadly categorized into policy constraint methods [49, 50], pessimistic value/policy regularization [51, 52], uncertainty estimation [53], importance sampling [54, 55], imitation learning [56, 57], and model-based methods [58, 59].
Theoretically, early offline RL studies relied on strong uniform data coverage assumptions [60, 61, 62, 63]. Recent works have relaxed these assumptions to partial coverage for tabular Markov Decision Processes (MDPs) [64, 65, 66, 67], linear MDPs [68, 69, 70], and general function approximation settings [71, 72, 73, 74].
Offline bandit learning has also been explored in multi-armed bandits (MAB) [64], contextual MABs [64, 68, 75], and neural contextual bandits [76, 77].
While our work leverages the pessimism principle and focuses on partial coverage settings, none of the aforementioned offline bandit or RL studies address the combinatorial action space, which is the central focus of our work. Conversely, recent work in the CMAB framework demonstrates that episodic tabular RL can be viewed as a special case of CMAB [13]. Building on this connection, our proposed framework can potentially extend to certain offline RL problems, offering a unified approach to tackle both combinatorial action spaces and offline learning.
1.1.3 Related Offline Learning Applications.
Cascading bandits, a classical online learning-to-rank framework, have been extensively studied in the literature [18, 26, 19, 78, 79, 80, 29, 81, 82]. Offline cascading bandits, on the other hand, focus primarily on reducing bias in learning settings [83, 84, 85, 86, 87]. Unlike these prior works, our study tackles the unbiased setting where data coverage is insufficient. Moreover, we are the first to provide a theoretically guaranteed solution using a CMAB-based approach.
LLM caching is a memory management technique aimed at mitigating memory footprints and access overhead during training and inference. Previous studies have investigated LLM caching at various levels, including attention-level (KV-cache) [88, 89, 90, 91], query-level [92, 34], and model/API-level [93, 94, 95]. Among these, the closest related work is the LLM cache bandit framework proposed by Zhu et al. [34]. However, their approach is ad hoc, whereas our CMAB-based framework systematically tackles the same problem and achieves improved results in both offline and online settings.
Influence Maximization (IM) was initially formulated as an algorithmic problem by Richardson and Domingos [96] and has since been studied using greedy approximation algorithms [97, 98]. The online IM problem has also received significant attention [99, 100, 101, 102, 103, 9]. In the offline IM domain, our work aligns closely with the optimization-from-samples (OPS) framework [104, 105, 106, 107], originally proposed by Balkanski et al. [104]. Specifically, our work falls under the subdomain of optimization-from-structured-samples (OPSS) [106, 107], where samples include detailed diffusion step information $(S_{0},...,S_{V-1})$ instead of only the final influence spread $\sigma(S_{0};G)$ in the standard OPS. Compared to Chen et al. [107], which selects the best seed set using empirical means, our approach employs a variance-adaptive pessimistic LCB, improving the suboptimal gap under relaxed assumptions.
@@ 2.1 Combinatorial Multi-armed Bandits with Probabilistically Triggered Arms
The original combinatorial multi-armed bandits problem with probabilistically triggered arms (CMAB-T) is an online learning game between a learner and the environment in $n$ rounds. We can specify a CMAB-T problem by a tuple :$\mathcal{I}\vcentcolon=([m],\mathbb{D},\mathcal{S},\mathbb{D}_{\text{trig}},R)$, where $[m]$ are base arms, $\mathcal{S}$ are the set of feasible combinatorial actions, $\mathbb{D}$ is the set of feasible distributions for the base arm outcomes, $\mathbb{D}_{\text{trig}}$ is the probabilistic triggering function, and $R$ is the reward function. The details of each component are described below:
@@
Combinatorial actions. At each round $t\in[n]$, the learner selects a combinatorial action $S_{t}\in\mathcal{S}$, where $\mathcal{S}$ is the set of feasible actions. Typically, $S_{t}$ is composed of a set of individual base arms $S\subseteq[m]$, which we refer to as a super arm. However, $S_{t}$ can be more general than the super arm, possibly continuous, such as resource allocations [108], which we will emphasize if needed.
Probabilistic arm triggering feedback. Motivated by the properties of real-world applications that will be introduced in detail in Section 4, we consider a feedback process that involves scenarios where each base arm in a super arm $S_{t}$ does not always reveal its outcome, even probabilistically. For example, a user might leave the system randomly at some point before examining the entire recommended list $S_{t}$, resulting in unobserved feedback for the unexamined items. To handle such probabilistic feedback, we assume that after the action $S_{t}$ is selected, the base arms in a random set $\tau_{t}\sim\mathbb{D}_{\text{trig}}(S_{t},\bm{X}_{t})$ are triggered depending on the outcome $\bm{X}_{t}$, where $\mathbb{D}_{\text{trig}}(S,\bm{X})$ is an unknown probabilistic distribution over the subsets $2^{[m]}$ given $S$ and $\bm{X}$. This means that the outcomes of the arms in $\tau_{t}$, i.e., $(X_{t,i})_{i\in\tau_{t}}$ are revealed as feedback to the learner, which could also be involved in determining the reward of action $S_{t}$ as we introduce later. To allow the algorithm to estimate the mean $\mu_{i}$ directly from samples, we assume the outcome does not depend on whether the arm $i$ is triggered, i.e., $\mathbb{E}_{\bm{X}\sim\mathbb{D}_{\text{arm}},\tau\sim\mathbb{D}_{\text{trig}}(S,\bm{X})}[X_{i}|i\in\tau]=\mathbb{E}_{\bm{X}\sim\mathbb{D}_{\text{arm}}}[X_{i}]$. We use $p_{i}^{\mathbb{D}_{\text{arm}},S}$ to denote the probability that base arm $i$ is triggered when the action is $S$ and the mean vector is $\bm{\mu}$.
Reward function. At the end of round $t\in[n]$, the learner receives a nonnegative reward $R_{t}=R(S_{t},\bm{X}_{t},\tau_{t})$, determined by action $S_{t}$, outcome $\bm{X}_{t}$, and triggered arm set $\tau_{t}$. Similarly to [9], we assume the expected reward to be $r(S_{t};\bm{\mu}_{t})\vcentcolon=\mathbb{E}[R(S_{t},\bm{X}_{t},\tau)]$, a function of the unknown mean vector $\bm{\mu}$, where the expectation is taken over the randomness of $\bm{X}_{t}$ and $\tau_{t}\sim\mathbb{D}_{\text{trig}}(S_{t},\bm{X}_{t})$.
Reward conditions. Owing to the nonlinearity of the reward and the combinatorial structure of the action, it is essential to give some conditions for the reward function to achieve any meaningful theoretical guarantee [9]. We consider the following conditions:
Condition 1 (Monotonicity, Wang and Chen [9]).
@@
Condition 2 (1-norm TPM Bounded Smoothness, Wang and Chen [9]).
@@
Condition 1 indicates the reward is monotonically increasing when the parameter $\bm{\mu}$ increases. In the learning to rank application (Section 4.1), for example, 1 means that when the purchase probability for each item increases, the total number of purchases for recommending any list of items also increases. Condition 2 bounds the reward smoothness/sensitivity. For Condition 2, the key feature is that the parameter change in each base arm $i$ is modulated by the triggering probability $p_{i}^{\bm{\mu},S}$. Intuitively, for base arm $i$ that is unlikely to be triggered/observed (small $p_{i}^{\bm{\mu},S}$), Condition 2 ensures that a large change in $\mu_{i}$ (due to insufficient observation) only causes a small change (multiplied by $p_{i}^{\bm{\mu},S}$) in reward, improving a $1/p_{\min}$ factor over Wang and Chen [9], where $p_{\min}$ is the minimum positive triggering probability. In learning to rank applications, for example, since users will never purchase an item if it is not examined, increasing or decreasing the purchase probability of an item that is unlikely to be examined (i.e., with small $p_{i}^{\bm{\mu},S}$) does not significantly affect the total number of purchases.
@@ 2.2 Offline Data Collection and Performance Metric
Offline dataset. Fix any CMAB-T problem $\mathcal{I}$ together with its underlying distribution $\mathbb{D}_{\text{arm}}$. We consider the offline learning setting, that is, the learner only has access to a dataset $\mathcal{D}$ consisting of $n$ feedback data $\mathcal{D}\vcentcolon=\left\{\left(S_{t},\tau_{t},(X_{t,i})_{i\in\tau_{t}}\right)\right\}_{t=1}^{n}$ collected a priori by an experimenter. Here, we assume the experimenter takes an unknown data collecting distribution $\mathbb{D}_{\mathcal{S}}$ over feasible actions $\mathcal{S}$, such that $S_{t}$ is generated i.i.d. from $S_{t}\sim\mathbb{D}_{\mathcal{S}}$ for any offline data $t\in[n]$. After $S_{t}$ is sampled, the environment generates outcome $\bm{X}_{t}\sim\mathbb{D}_{\text{arm}}$. Then $\tau_{t}\sim\mathbb{D}_{\text{trig}}(S_{t},\bm{X}_{t})$ are triggered, whose outcome are recorded as $(X_{t,i})_{i\in\tau_{t}}$. To this end, we use $p_{i}^{\mathbb{D}_{\text{arm}},\mathbb{D}_{\mathcal{S}}}$ to denote the data triggering probability, i.e., $p_{i}^{\mathbb{D}_{\text{arm}},\mathbb{D}_{\mathcal{S}}}=\mathbb{E}_{S\sim\mathbb{D}_{\mathcal{S}},\bm{X}\sim\mathbb{D}_{\text{arm}},\tau\sim\mathbb{D}_{\text{trig}}(S,\bm{X})}[\mathbb{I}\left\{i\in\tau\right\}]$, which indicates the frequency of observing arm $i\in[m]$. To this end, we use $\mathbb{D}_{\text{joint}}\vcentcolon=(\mathbb{D}_{\mathcal{S}},\mathbb{D}_{\text{arm}},\mathbb{D}_{\text{trig}})$ to denote the joint distribution considering all possible randomness from the data collection $\mathbb{D}_{\mathcal{S}}$, the random base arm outcome $\mathbb{D}_{\text{arm}}$, and the probabilistic triggering $\mathbb{D}_{\text{trig}}$.
Approximation oracle and $\alpha$ approximate suboptimality gap. The goal of the offline learning problem for CMAB-T is to identify the optimal combinatorial action that maximizes the expected reward. Correspondingly, the performance of an offline learning algorithm $A$ is measured by the suboptimality-gap, defined as the difference in the expected reward between the optimal action $S^{*}\vcentcolon=\operatorname*{argmax}_{S^{\prime}\in\mathcal{S}}r(S^{\prime};\bm{\mu})$ and the action $\hat{S}$ chosen by algorithm $A$ with dataset $\mathcal{D}$ as input. For many reward functions, it is NP-hard to compute the exact $S^{*}$ even when $\bm{\mu}$ is known, so similar to [109, 9, 29], we assume that algorithm $A$ has access to an offline $\alpha$-approximation $\mathrm{ORACLE}$, which takes any mean vector $\bm{\mu}\in[0,1]^{m}$ as input, and outputs an $\alpha$-approximate solution $S\in\mathcal{S}$, i.e., $S=\texttt{ORACLE}(\bm{\mu})$ satisfies

### v2

@@ Abstract
The combinatorial multi-armed bandit (CMAB) is a fundamental sequential decision-making framework, extensively studied over the past decade. However, existing work primarily focuses on the online setting, overlooking the substantial costs of online interactions and the readily available offline datasets. To overcome these limitations, we introduce Off-CMAB, the first offline learning framework for CMAB. Central to our framework is the combinatorial lower confidence bound (CLCB) algorithm, which combines pessimistic reward estimations with combinatorial solvers. To characterize the quality of offline datasets, we propose two novel data coverage conditions and prove that, under these conditions, CLCB achieves a near-optimal suboptimality gap, matching the theoretical lower bound up to a logarithmic factor. We validate Off-CMAB through practical applications, including learning to rank, large language model (LLM) caching, and social influence maximization, showing its ability to handle nonlinear reward functions, general feedback models, and out-of-distribution action samples that exclude optimal or even feasible actions. Extensive experiments on synthetic and real-world datasets for these applications further highlight the superior performance of CLCB.
ABCD, ICML
@@ 1 Introduction
Combinatorial multi-armed bandit (CMAB) is a fundamental sequential decision-making framework, designed to tackle challenges in combinatorial action spaces. Over the past decade, CMAB has been extensively studied (Cesa-Bianchi & Lugosi, 2012; Bubeck et al., 2012; Audibert et al., 2014; Neu, 2015; Gai et al., 2012; Kveton et al., 2015c; Combes et al., 2015; Chen et al., 2016; Wang & Chen, 2017; Merlis & Mannor, 2019; Saha & Gopalan, 2019; Zimmert et al., 2019; Liu et al., 2024b; Qin et al., 2014; Liu et al., 2023a; Choi et al., 2024; Hwang et al., 2023), driving advancements in real-world applications like recommendation systems (Kveton et al., 2015a; Li et al., 2016; Lattimore et al., 2018; Agrawal et al., 2019), healthcare (Lin & Bouneffouf, 2022; Verma et al., 2023; Bouneffouf et al., 2020), and cyber-physical systems (György et al., 2007; Kveton et al., 2015b; Li et al., 2019; Liu et al., 2023b).
@@
Most success stories of CMAB have emerged within the realm of online CMAB [^1] See Appendix A for a comprehensive discussion on the related works., which relies on active data collection through online exploration. While effective in certain scenarios, this framework faces two major limitations. On one hand, online exploration becomes impractical when it incurs prohibitive costs or raises ethical and safety concerns. On the other hand, they neglect offline datasets that are often readily available at little or no cost.
For instance, in healthcare systems (Liu et al., 2020), recommending optimal combinations of medical treatments—such as drugs, surgical procedures, and radiation therapy—requires extreme caution. Experimenting directly on patients is ethically and practically infeasible. Instead, leveraging pre-collected datasets of prior treatments can help to make informed decisions while ensuring patient safety. Similar happens for recommendation systems (Chen et al., 2023) and autonomous driving (Kiran et al., 2020), offline datasets such as user click histories and human driving logs are ubiquitous. Leveraging these offline datasets can guide learning agents to identify optimal policies while avoiding the significant costs associated with online exploration—such as degrading user experience or risking car accidents. For more examples, see Appendix B for details.
@@
To address the limitations of online CMAB, we propose the first offline learning framework for CMAB (Off-CMAB), where we leverage a pre-collected dataset consisting of $n$ samples of combinatorial actions and their corresponding feedback data. Our framework handles rewards that are nonlinear functions of the chosen super arms and considers probabilistic feedback models that generalize the standard semi-bandit feedback model (Gai et al., 2010; Chen et al., 2013; Kveton et al., 2015c), supporting a wide range of applications such as learning to rank (Liu et al., 2009), large language model (LLM) caching (Zhu et al., 2023), and influence maximization (Kempe et al., 2003a). The objective is to identify a combinatorial action that minimizes the suboptimal gap, defined as the reward difference between the optimal action and the identified action.
The key challenge of Off-CMAB lies in the absence of access to an online environment, which inherently limits the number of data samples available for each action. Furthermore, this problem becomes even more challenging with the combinatorially large action space, which complicates the search for optimal solutions, and the potential presence of out-of-distribution (OOD) samples, where the dataset may exclude optimal or even feasible actions. To tackle these challenges, this work makes progress in answering the following two open questions:
@@
Practical Applications: We show the practicality of Off-CMAB by fitting real-world problems into our framework and applying CLCB to solve them, including (1) learning to rank, (2) LLM caching, and (3) social influence maximization (IM). For the LLM cache problem, beyond directly fitting it into our framework, we improve existing results by addressing full-feedback arms, extending our approach to the online LLM setting with similar improvements. For social IM, our framework handles nuanced node-level feedback by constructing base-arm LCBs via intermediate UCB/LCBs, with additional refinements using variance-adaptive confidence intervals for improved performance.
@@ 2.1 Combinatorial Multi-armed Bandits with Probabilistically Triggered Arms
The original combinatorial multi-armed bandits problem with probabilistically triggered arms (CMAB-T) is an online learning game between a learner and the environment in $n$ rounds. We can specify a CMAB-T problem by a tuple :$\mathcal{I}\vcentcolon=([m],\mathbb{D},\mathcal{S},\mathbb{D}_{\text{trig}},R)$, where $[m]$ are base arms, $\mathcal{S}$ is the set of feasible combinatorial actions, $\mathbb{D}$ is the set of feasible distributions for the base arm outcomes, $\mathbb{D}_{\text{trig}}$ is the probabilistic triggering function, and $R$ is the reward function. The details of each component are described below:
@@
Combinatorial actions. At each round $t\in[n]$, the learner selects a combinatorial action $S_{t}\in\mathcal{S}$, where $\mathcal{S}$ is the set of feasible actions. Typically, $S_{t}$ is a set of individual base arms $S\subseteq[m]$, which we refer to as a super arm. However, $S_{t}$ can be more general than the super arm, e.g., continuous arms are useful for applying CMAB to resource allocation (Zuo & Joe-Wong, 2021), which we emphasize as needed.
Probabilistic arm triggering feedback. Motivated by the properties of real-world applications that will be introduced in detail in Section 4, we consider a feedback process that involves scenarios where each base arm in a super arm $S_{t}$ does not always reveal its outcome, even probabilistically. For example, a user might leave the system randomly at some point before examining the entire recommended list $S_{t}$, resulting in unobserved feedback for the unexamined items. To handle such probabilistic feedback, we assume that after the action $S_{t}$ is selected, the base arms in a random set $\tau_{t}\sim\mathbb{D}_{\text{trig}}(S_{t},\bm{X}_{t})$ are triggered depending on the outcome $\bm{X}_{t}$, where $\mathbb{D}_{\text{trig}}(S,\bm{X})$ is an unknown probabilistic distribution over the subsets $2^{[m]}$ given $S$ and $\bm{X}$. This means that the outcomes of the arms in $\tau_{t}$, i.e., $(X_{t,i})_{i\in\tau_{t}}$, are revealed as feedback to the learner, which could also be involved in determining the reward of action $S_{t}$ as we introduce later. To allow the algorithm to estimate the mean $\mu_{i}$ directly from samples, we assume the outcome does not depend on whether the arm $i$ is triggered, i.e., $\mathbb{E}_{\bm{X}\sim\mathbb{D}_{\text{arm}},\tau\sim\mathbb{D}_{\text{trig}}(S,\bm{X})}[X_{i}|i\in\tau]=\mathbb{E}_{\bm{X}\sim\mathbb{D}_{\text{arm}}}[X_{i}]$. We use $p_{i}^{\mathbb{D}_{\text{arm}},S}$ to denote the probability that base arm $i$ is triggered when the action is $S$ and the mean vector is $\bm{\mu}$.
Reward function. At the end of round $t\in[n]$, the learner receives a nonnegative reward $R_{t}=R(S_{t},\bm{X}_{t},\tau_{t})$, determined by action $S_{t}$, outcome $\bm{X}_{t}$, and triggered arm set $\tau_{t}$. Similarly to (Wang & Chen, 2017), we assume the expected reward to be $r(S_{t};\bm{\mu}_{t})\vcentcolon=\mathbb{E}[R(S_{t},\bm{X}_{t},\tau)]$, a function of the unknown mean vector $\bm{\mu}$, where the expectation is taken over the randomness of $\bm{X}_{t}$ and $\tau_{t}\sim\mathbb{D}_{\text{trig}}(S_{t},\bm{X}_{t})$.
Reward conditions. Owing to the nonlinearity of the reward and the combinatorial structure of the action, it is essential to give some conditions for the reward function to achieve any meaningful theoretical guarantee (Wang & Chen, 2017). We consider the following conditions:
Condition 1 (Monotonicity, Wang & Chen (2017)).
@@
Condition 2 (1-norm TPM Bounded Smoothness, Wang & Chen (2017)).
@@
Condition 1 indicates the reward is monotonically increasing when the parameter $\bm{\mu}$ increases. Condition 2 bounds the reward smoothness/sensitivity, i.e., the amount of the reward change caused by the parameter change from $\bm{\mu}$ to $\bm{\mu}^{\prime}$. In the learning to rank (Section 4.1), for example, these conditions upper bounds the difference in total number of purchases when the purchase probability for the items changes from $\bm{\mu}$ to $\bm{\mu}^{\prime}$. For Condition 2, the key feature is that the parameter change in each base arm $i$ is modulated by the triggering probability $p_{i}^{\bm{\mu},S}$, saving a $p_{\min}$ factor in (Chen et al., 2016) where $p_{\min}$ is the minimum positive triggering probability. Intuitively, for base arm $i$ that is unlikely to be triggered/observed (small $p_{i}^{\bm{\mu},S}$), Condition 2 ensures that a large change in $\mu_{i}$ (due to insufficient observation) only causes a small change (multiplied by $p_{i}^{\bm{\mu},S}$) in reward, saving a $p_{\min}$ factor in (Wang & Chen, 2017) where $p_{\min}$ is the minimum positive triggering probability. In learning to rank application, for example, since users will never purchase an item if it is not examined, increasing or decreasing the purchase probability of an item that is unlikely to be examined (i.e., with small $p_{i}^{\bm{\mu},S}$) does not significantly affect the total number of purchases.
@@ 2.2 Offline Data Collection and Performance Metric
Offline dataset. Fix any CMAB-T problem $\mathcal{I}$ together with its underlying distribution $\mathbb{D}_{\text{arm}}$. We consider the offline learning setting, that is, the learner only has access to a dataset $\mathcal{D}$ consisting of $n$ feedback data $\mathcal{D}\vcentcolon=\left\{\left(S_{t},\tau_{t},(X_{t,i})_{i\in\tau_{t}}\right)\right\}_{t=1}^{n}$ collected a priori by an experimenter. Here, we assume the experimenter takes an unknown data collecting distribution $\mathbb{D}_{\mathcal{S}}$ over feasible actions $\mathcal{S}$, such that $S_{t}$ is generated i.i.d. from $S_{t}\sim\mathbb{D}_{\mathcal{S}}$ for any offline data $t\in[n]$. After $S_{t}$ is sampled, the environment generates outcome $\bm{X}_{t}\sim\mathbb{D}_{\text{arm}}$. Then $\tau_{t}\sim\mathbb{D}_{\text{trig}}(S_{t},\bm{X}_{t})$ are triggered, whose outcome are recorded as $(X_{t,i})_{i\in\tau_{t}}$. To this end, we use $p_{i}^{\mathbb{D}_{\text{arm}},\mathbb{D}_{\mathcal{S}}}$ to denote the data triggering probability, i.e., $p_{i}^{\mathbb{D}_{\text{arm}},\mathbb{D}_{\mathcal{S}}}=\mathbb{E}_{S\sim\mathbb{D}_{\mathcal{S}},\bm{X}\sim\mathbb{D}_{\text{arm}},\tau\sim\mathbb{D}_{\text{trig}}(S,\bm{X})}[\mathbb{I}\left\{i\in\tau\right\}]$, which indicates the frequency of observing arm $i\in[m]$.
Approximation oracle and $\alpha$ approximate suboptimality gap. The goal of the offline learning problem for CMAB-T is to identify the optimal combinatorial action that maximizes the expected reward. Correspondingly, the performance of an offline learning algorithm $A$ is measured by the suboptimality-gap, defined as the difference in the expected reward between the optimal action $S^{*}\vcentcolon=\operatorname*{argmax}_{S^{\prime}\in\mathcal{S}}r(S^{\prime};\bm{\mu})$ and the action $\hat{S}$ chosen by algorithm $A$ with dataset $\mathcal{D}$ as input. For many reward functions, it is NP-hard to compute the exact $S^{*}$ even when $\bm{\mu}$ is known, so similar to (Chen et al., 2013; Wang & Chen, 2017; Liu et al., 2022, 2024a), we assume that algorithm $A$ has access to an offline $\alpha$-approximation $\mathrm{ORACLE}$, which takes any mean vector $\bm{\mu}\in[0,1]^{m}$ as input, and outputs an $\alpha$-approximate solution $S\in\mathcal{S}$, i.e., $S=\texttt{ORACLE}(\bm{\mu})$ satisfies